You want to make 45 ml. Of a 35% alcohol solution. You call the pharmacy and they tell you that all they have is 60% alcohol. How much of the 60% alcohol solution should you ask for and how much water should you add?
Answer by ptaylor(2198)

Let x=amount of 60% alcohol that I need
And 45-x=amount of water that I should add
Now we know that the amount of pure alcohol that exists before the mixture takes place(0.60x) has to equal the amount of pure alcohol that exists after the mixture takes place(0.35*45)
Sooooo
0.60x=0.35*45=15.75
x=(15.75)/0.60=26.25ml---- amount of 60% alcohol needed
45-x=45-26.25=18.75ml-----amount of water needed
CK
0.60*26.25=???15.75
15.75=15.75
Hope this helps---ptaylor